Find all second order partial derivatives of f(x, y) = (x² +3y)
Answer #1

f(x,y) = (x+3y)? fas 36**+ 3y) (2x). 6x (x + za)? for = 6(x (2(x²+34*) (2x)) + pe ²+3ye). ] 6[ ux(x++3yo) + (x + 3yo)?J 6[(x